# Partial Law Invariance and Risk Measures

### v4

@@ 3.3 Two constructions
where $\widehat{\rho}^{\beta}$ is the $\beta$-tail risk measure generated by $\widehat{\rho}$. See Section C.3 for a proof of this claim.
4 Multiple sources
A natural corollary of Theorem 1 characterizes all Fatou $\mathfrak{G}$-law invariant coherent risk measures.
Theorem 2.
$\rho(X)=\sup_{\mu\in\mathcal{S}}\mathbb{E}^{\mu}[X],\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}$
Appendix D contains an example of a $\mathfrak{G}$-law-invariant coherent risk measure given by $\rho(X)=\sup_{\mathcal{G}\in\mathfrak{G}}\rho_{\mathcal{G}}(\mathbb{E}[X|\mathcal{G}])$, where each $\rho_{\mathcal{G}}$ is law invariant and coherent, along with its representation in the form of Theorem 2.
  1. (a)
  2. (b)
It is clear that both (a) and (b) are stronger than $\mathfrak{G}$-law invariance, and moreover, (b) is stronger than (a). In Example 6 below, we illustrate a situation in which (b) is not desirable, whereas (a) may be reasonable.
Example 6.
$\rho(X)=\sup_{\lambda\in[0,1]}\mathrm{ES}^{\mu_{\lambda}}_{\alpha}(X),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
Our current formulation of $\mathfrak{G}$-law invariance has an advantage over property (a), that is, the flexibility for modeling the situation where different sources are associated with different levels of model uncertainty (ambiguity), illustrated by Example 7 below.
Example 7.
$\rho(X)=\sup_{\mu\in\mathcal{M}_{1}}\left\{\mathbb{E}^{\mu}[\ell(X)]-\frac{1}{\theta_{i}}R(\mu,\mathbb{P})\right\}=\frac{1}{\theta_{i}}\log\mathbb{E}\left[e^{\theta_{i}\ell(X)}\right],\leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}(\mathcal{G}_{i}),$ (10)
In Example 7, even if model uncertainty exists in each source $\mathcal{G}_{1},\dots,\mathcal{G}_{n}$, the risk measure is law invariant on each subspace as a result of robustification, discussed by Hansen and Sargent (2001).
Example 8.
$\mathrm{ER}_{\beta}(X)=\frac{1}{\beta}\log\mathbb{E}\left[e^{\beta X}\right],\text{\leavevmode\nobreak\ for\leavevmode\nobreak\ $\beta>0$},\text{\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ $\mathrm{ER}_{0}=\mathbb{E}$}.$
5 Extensions
In this section, we present three extensions of partially law-invaraint coherent risk measures. First, in Section 5.1, we introduce a technical tool called coherent adjustments, which admit interesting mathematical properties. Coherent adjustments allow us to reformulate Theorem 1 and serve as the foundation for the subsequent two extensions. Second, in Section 5.2, we strengthen partial law invariance to a property, which still generalizes law invariance, that allows us to extend the representation of Kusuoka (2001). Finally, in Section 5.3, we extend some of the results regarding partially law-invariant coherent risk measures to partially law-invariant convex risk measures.
5.1 Coherent adjustments and an alternative representation
Definition 1.
Let $\mu\in\mathcal{M}_{1}(\mathcal{G})$, from the above properties, it is straightforward that the supremum of $\mu$-coherent adjustments is again a $\mu$-coherent adjustment. A simple example of a Fatou $\mu$-coherent adjustment is $\tau=\mathbb{E}^{\nu}-\mathbb{E}^{\mu}$ for any $\nu\in\mathcal{E}(\mu)$, which is easy to check. In Section E.1, Proposition 17, shows that, indeed, the supremum of the form $\tau=\mathbb{E}^{\nu}-\mathbb{E}^{\mu}$ contains all possible Fatou $\mu$-coherent adjustments.
Remark 2.
The name “$\mu$-coherent adjustment” is explained by Proposition 8.
Proposition 8.
  1. (i)
  2. (ii)
If we assume additivity on $L^{\infty}(\mathcal{G})$ and $\mathcal{G}$-law invariance for a Fatou coherent risk measure $\rho$, then $\rho$ has a form $\mathbb{E}^{\mathbb{P}}+\tau$ for $\tau\in\mathrm{CA}(\mathbb{P}|_{\mathcal{G}})$ that is similar to Proposition 8 (ii); see Section E.2.
For some $\mathcal{R}\subseteq\mathcal{M}_{1}(\mathcal{G})$ and CA-assignment $(\tau_{\mu})_{\mu\in\mathcal{R}}$, by Proposition 8, the following risk measure
$\rho(X)=\sup_{\mu\in\mathcal{R}}\left\{\mathbb{E}^{\mu}[X]+\tau_{\mu}(X)\right\},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty},$ (11)
is coherent, as it is the supremum of coherent risk measures. Theorem 3 below shows that this mapping is Fatou continuous and that all Fatou coherent risk measures have this representation.
To explain why $\mu$-coherent adjustments help us extend a coherent risk measure on $L^{\infty}(\mathcal{G})$ to $L^{\infty}$, let $\tilde{\rho}:L^{\infty}(\mathcal{G})\to\mathbb{R}$ be a law-invariant coherent risk measure. For a convex $\mathcal{R}\subseteq\mathcal{M}_{1}(\mathcal{G})$ that is $w(\mathcal{M}(\mathcal{G}),L^{\infty}(\mathcal{G}))$-dense in $\mathcal{S}_{\tilde{\rho}}$ and a CA-assignment $(\tau_{\mu})_{\mu\in\mathcal{R}}$, let $\rho$ be given by (11). It is easy to see that $\rho|_{L^{\infty}(\mathcal{G})}=\tilde{\rho}$, meaning that it is $\mathcal{G}$-law invariant. In the special case that $\tau_{\mu}=0$ for all $\mu\in\mathcal{R}$, this extension technique precisely gives the class studied in Proposition 6, because $\widehat{\mathcal{S}}_{\rho}=\widehat{\mathcal{S}}_{\tilde{\rho}}\subseteq L^{1}(\mathcal{G})$.
Theorem 3.
$\rho(X)=\sup_{\mu\in\mathcal{R}}\left\{\mathbb{E}^{\mu}[X]+\tau_{\mu}(X)\right\},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$ (12)
Remark 3.
There is a straightforward way to switch between representations (4) and (12) for a Fatou and $\mathcal{G}$-law-invariant coherent risk measure. Given a convex set $\mathcal{S}\subseteq\mathcal{M}_{1}(\mathcal{G})$ that represents $\rho$ in Theorem 1, define $\mathcal{R}=\mathcal{S}^{\mathcal{G}}$, which is almost $\mathcal{G}$-law invariant. For each $\mu\in\mathcal{R}$, define $\mathcal{S}(\mu)=\mathcal{E}(\mu)\cap\mathcal{S}$ and $\tau_{\mu}(X)=\sup_{\nu\in\mathcal{S}(\mu)}\{\mathbb{E}^{\nu}[X]-\mathbb{E}^{\mu}[X]\}$. These can be used in representation (12). Conversely, let an almost $\mathcal{G}$-law invariant convex set $\mathcal{R}\subseteq\mathcal{M}_{1}(\mathcal{G})$ and a CA-assignment $(\tau_{\mu})_{\mu\in\mathcal{R}}$ representing $\rho$ in Theorem 3 be given. For $\mu\in\mathcal{R}$, let $\mathcal{S}_{\tau_{\mu}}$ be the $w(\mathcal{M},L^{\infty})$-closed and convex set in Proposition 17 (see Section E.1) that represents $\tau_{\mu}$. It is easy to show that $\mathcal{S}=\bigcup_{\mu\in\mathcal{R}}\mathcal{S}_{\tau_{\mu}}$ is convex and $S^{\mathcal{G}}=\mathcal{R}$. We can use $\mathcal{S}$ in representation (4). The above connection clarifies the construction in (5).
Uniqueness of representation (12) under a stronger notion of continuity is addressed in Section E.3. Representation (12) has little resemblance to the classical representation result of Kusuoka (2001) for law-invariant coherent risk measures. The following section addresses this gap by further imposing a necessary and sufficient condition for a similar representation along that direction.
5.2 Strong partial law invariance and Kusuoka-type representation
Next, we introduce a new property, which is stronger than partial law invariance and is necessary and sufficient to reduce (12) to a Kusuoka-type representation; see Proposition 11 in Appendix A.
$\rho(Z+X)=\rho(Z+Y).$
Lemma 1.
The mapping $\rho$ is strongly $\mathcal{G}$-law invariant if and only if for all $X\in L^{\infty}$ and $Y\in L^{\infty}(\mathcal{G})$ with $Y\overset{\mathrm{d}}{=}_{\mathbb{P}}\mathbb{E}[X|\mathcal{G}]$ we have
$\rho(X)=\rho(X-\mathbb{E}[X|\mathcal{G}]+Y).$
An example of strongly $\mathcal{G}$-law invariant risk measure $\rho$ is $\rho(X)=\tilde{\rho}(\mathbb{E}[X|\mathcal{G}])$ given in (7), where $\tilde{\rho}:L^{\infty}(\mathcal{G})\to\mathbb{R}$ is law invariant. Examples of functionals satisfying $\mathcal{G}$-law invariance but not strong $\mathcal{G}$-law invariance will be provided in Section E.5.
$Q([0,t))=\int_{[0,t)}(1-s)\leavevmode\nobreak\ \mathrm{d}q_{D_{\mu}}(s),\text{\leavevmode\nobreak\ \leavevmode\nobreak\ for\leavevmode\nobreak\ \leavevmode\nobreak\ }t\in[0,1).$ (13)
[…]
$\sup_{\mu\in\mathcal{M}_{Q}}\mathbb{E}^{\mu}[X]=\int_{[0,1)}\mathrm{ES}_{\alpha}(\mathbb{E}[X|\mathcal{G}])\leavevmode\nobreak\ Q(\mathrm{d}\alpha);$ (14)
Theorem 4.
Let $\rho:L^{\infty}\to\mathbb{R}$, the following are equivalent.
  1. (i)
  2. (ii)
    $\rho(X)=\sup_{Q\in\mathcal{D}}\left\{\int_{[0,1)}\mathrm{ES}_{\alpha}\left(\mathbb{E}[X|\mathcal{G}]\right)\leavevmode\nobreak\ Q(\mathrm{d}\alpha)+\tau_{Q}(X)\right\}.$ (15)
$\rho(X)=\sup_{Q\in\mathcal{D}}\int_{[0,1)}\mathrm{ES}_{\alpha}(X)\leavevmode\nobreak\ Q(\mathrm{d}\alpha),\hskip 3.0ptX\in L^{\infty},$
because in this case $\mathbb{E}[X|\mathcal{G}]=X$, $\tau_{Q}(X)=0$, and strong $\mathcal{G}$-law invariance is law invariance. A simple example is to choose $\tau_{Q}=0$ for all $Q\in\mathcal{D}$, which yields the risk measure $\rho:X\mapsto\tilde{\rho}(\mathbb{E}[X|\mathcal{G}])$ studied in Proposition 6.
5.3 Convex risk measures
This section discusses convex risk measures and their relation to the variational preferences model of Maccheroni et al. (2006). We provide representations for partially law-invariant convex risk measures akin to Theorem 3. A risk measure $\rho:L^{\infty}\to\mathbb{R}$ is Fatou and convex if and only if there exists $\alpha:\mathcal{M}_{1}\to(-\infty,\infty]$ such that
$\rho(X)=\sup_{\mu\in\mathcal{M}_{1}}(\mathbb{E}^{\mu}[X]-\alpha(\mu)),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty};$
Similar to CA-assignments, we say that $(\tau_{\mu})_{\mu\in\mathcal{M}_{1}(\mathcal{G})}$ is a CoA-assignment if for all $\mu\in\mathcal{M}_{1}(\mathcal{G})$, $\tau_{\mu}$ is a Fatou $\mu$-convex adjustment and the mapping $\mu\mapsto\tau_{\mu}(0)$ is concave. The following result is the convex version of Theorem 3.
Theorem 5.
$\rho(X)=\sup_{\mu\in\mathcal{M}_{1}(\mathcal{G})}\left\{\mathbb{E}^{\mu}[X]+\tau_{\mu}(X)\right\}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
The risk measures ER in Example 8 are the most common examples of convex risk measures. We next discuss a few $\mathcal{G}$-law-invariant versions of ER, which shed some light on different ways to construct partially law-invariant convex risk measures. In the examples below, partial law invariance is not necessarily associated with the interpretation of model uncertainty.
$\mathrm{ER}_{\beta}(X|\mathcal{G})=\frac{1}{\beta}\log\mathbb{E}\left[e^{\beta X}|\mathcal{G}\right],\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
${\mathrm{ER}}^{\mathcal{G}}_{\beta}(X)=\mathbb{E}[\mathrm{ER}_{\beta}(X|\mathcal{G})]=\mathbb{E}\left[\frac{1}{\beta}\log\mathbb{E}\left[e^{\beta X}|\mathcal{G}\right]\right],\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
The second is to compute the ER of the conditional mean in the form of (7), that is,
$\widetilde{\mathrm{ER}}^{\mathcal{G}}_{\beta}(X)=\mathrm{ER}_{\beta}(\mathbb{E}[X|\mathcal{G}])=\frac{1}{\beta}\log\mathbb{E}\left[e^{\beta\mathbb{E}[X|\mathcal{G}]}\right],\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
$\mathrm{ER}_{\beta}(\mathrm{ER}_{\beta}(X|\mathcal{G}))=\frac{1}{\beta}\log\mathbb{E}\left[\exp\left(\beta\frac{1}{\beta}\log\mathbb{E}\left[e^{\beta X}|\mathcal{G}\right]\right)\right]=\mathrm{ER}_{\beta}(X),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
${\mathrm{ER}}^{\mathcal{G}}_{\beta}(X)=\mathbb{E}[X]\mbox{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ }\widetilde{\mathrm{ER}}^{\mathcal{G}}_{\beta}(X)=\mathrm{ER}_{\beta}(X);$
if $Y$ is independent of $\mathcal{G}$, then
${\mathrm{ER}}^{\mathcal{G}}_{\beta}(Y)=\mathrm{ER}_{\beta}(Y)\mbox{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ }\widetilde{\mathrm{ER}}^{\mathcal{G}}_{\beta}(Y)=\mathbb{E}[Y].$
6 Partially law-invariant ES and numerical illustrations
$\overline{\mathrm{ES}}^{\mathcal{Q}}_{\alpha}(X)=\sup_{\mu\in\mathcal{Q}}\mathrm{ES}^{\mu}_{\alpha}(X),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty},$
where $\mathcal{Q}\subseteq\mathcal{E}(\mathbb{P})$ and satisfies some regularity (to be made rigorous later). Afterward, we numerically implement the section’s results and see how decisions in a simple investment vary as we change the parameters in the risk measure.
We focus on ES here for two reasons. First, ES is the standard regulatory risk measure in the banking sector (BCBS (2019)), and hence the most relevant for financial applications. Second, ES has special advantages in optimization (Rockafellar and Uryasev (2002)), and the partially law-invariant version of ES admits tractable formulas (Proposition 9 below).
Definition 2.
  1. (i)
    For all $X\in\mathcal{X}$, $X\leq Y$;
  2. (ii)
    if $Z\in L^{\infty}$ and $X\leq Z$ for all $X\in\mathcal{X}$, then $Y\leq Z$.
Definition 3.
Given certain conditions on $\mathcal{Q}\subseteq\mathcal{E}(\mathbb{P})$, the $\mathcal{G}$-law invariant risk measure $\sup_{\mu\in\mathcal{Q}}\mathrm{ES}^{\mu}$ mentioned in Section 2.2 can be determined via a minimization formula, similar to the formulation of ES by Rockafellar and Uryasev (2002). Below, a coherent conditional risk measure is a mapping from $L^{\infty}$ to $L^{\infty}(\mathcal{G})$ satisfying the for axioms of coherence but with constants replaced by elements of $L^{\infty}(\mathcal{G})$; see Detlefsen and Scandolo (2005).
Proposition 9.
$\sup_{\mu\in\mathcal{Q}}\mathrm{ES}^{\mu}_{\alpha}(X)=\min_{x\in\mathbb{R}}\left(x+\frac{1}{1-\alpha}\mathbb{E}\left[\rho_{\mathcal{G}}^{\mathcal{Q}}\left((X-x)^{+}\right)\right]\right),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty},$
$\rho_{\mathcal{G}}^{\mathcal{Q}}(X)=\operatorname*{ess\text{-}sup}_{\mu\in\mathcal{Q}}\mathbb{E}^{\mu}[X|\mathcal{G}],\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ X\in L^{\infty}.$
Moreover, the set $\mathcal{Q}_{\beta}=\left\{\mu\in\mathcal{E}(\mathbb{P}):\leavevmode\nobreak\ D_{\mu}\leq 1/(1-\beta)\right\}$ for $\beta\in[0,1)$ is $w(\mathcal{M},L^{\infty})$-compact and $\mathcal{G}$-convex.
Denote by $\mathrm{ES}_{\beta}(X|\mathcal{G})=\rho_{\mathcal{G}}^{\mathcal{Q}_{\beta}}(X)$, $X\in L^{\infty}$, which is the conditional version of ES in Detlefsen and Scandolo (2005). The benefit of using $\mathcal{Q}_{\beta}$ is that under tractable assumptions of the conditional distribution of $X$ given $\mathcal{G}$, $\mathbb{E}[\mathrm{ES}_{\beta}((X-x)^{+}|\mathcal{G})]$ can be computed using conditional distributions.
Next, we illustrate with a Gaussian example in the context of investment. Let $(\Omega,\mathcal{F})=(\mathbb{R}^{2},\mathcal{B}(\mathbb{R}^{2}))$ and $\mathbb{P}$ be multivariate Gaussian with mean $m\in\mathbb{R}^{2}$ and covariance matrix $\Sigma\in\mathbb{R}^{2\times 2}$, where $\Sigma_{ij}=(c+(1-c)\mathds{1}_{\{i=j\}})\sigma_{i}\sigma_{j}$ for $i,j\in\{1,2\}$. Let $X_{1},X_{2}$ be the random variables given by the projections. We interpret $X_{1},X_{2}$ as the losses (as a percentage) of two positions on a market and $\mathbb{P}$ as their estimated bivariate distribution. Define $\mathcal{G}=\sigma(X_{1})$, the goal is to use the risk measure $\rho_{\beta}=\sup_{\mu\in\mathcal{Q}^{\beta}}\mathrm{ES}^{\mu}_{\alpha}$ to measure the risk between different portfolios given by $\pi_{1}X_{1}+\pi_{2}X_{2}$, where $(\pi_{1},\pi_{2})$ is a vector in the unit simplex in $\mathbb{R}^{2}$.
In Appendix F, for $\beta\in[0,1)$, we derive an integral formula for
@@ 6 Partially law-invariant ES and numerical illustrations
which we can compute via numerical integration. Therefore, by Proposition 9,
$\overline{\mathrm{ES}}_{\alpha}^{\mathcal{Q}_{\beta}}\left(\pi_{1}X_{1}+\pi_{2}X_{2}\right)=\min_{x\in\mathbb{R}}\left(x+\frac{1}{1-\alpha}f_{\beta}(\pi_{1},\pi_{2},x)\right),$
which we can compute via convex optimization.
Figures 1 and 2 display plots depicting the risk, calculated using $\rho_{\beta}$, associated with various portfolios. The horizontal axes in all plots represent the weight allocated to $X_{1}$. In all cases, $\rho_{\beta}(X_{1})=\mathrm{ES}_{\alpha}(X_{1})$, explaining why all the lines agree when $\pi_{1}=1$.
Figure 1 aligns with our theoretical expectations. Specifically, when $m_{2}=0$ (plot 1 in Figure 1), an agent who assesses their risk using $\beta=0$ opts for a balanced investment in $X_{1}$ and $X_{2}$. This conclusion arises since the risk measure $\rho_{0}$ coincides with $\mathrm{ES}_{\alpha}$, and the agent attains the lowest risk through total diversification. As $\beta$ increases, the optimally risk-minimized portfolio progressively shifts towards an investment completely in $X_{1}$. This happens because the agent is becoming increasingly fearful that the distribution of $X_{2}$ is misspecified. The additional plots within Figure 1 explore the scenarios when $m_{2}=-0.05,-0.1$, indicating instances when $X_{2}$ is presumed to yield a positive average return. The observed trend mirrors the behavior noted in the scenario where $m_{2}=0$, albeit with reduced intensity.
Figure 1: Varying $\beta$, $m_{1}=0$, $\sigma_{1}=0.1$, $\sigma_{2}=0.1$, $c=0.5$, $\alpha=0.95$; the red dotted curve connects the optimizers for different $\beta$
Figure 2 fixes $\beta=0.95$ but varies $\sigma_{2}$. In all plots, an agent who assesses their risk using $\rho_{0.95}$ invests fully into $X_{2}$ when the standard deviation of $X_{2}$ is low enough. The agent fears model misspecification, with the parameter $\beta$ quantifying this fear. Since $\beta=0.95$, the agent heavily suspects that $X_{2}$ is misspecified. However, when $\sigma_{2}$ is very small, because the distributions in “the $\beta$-neighbourhood” of the estimated distribution of $X_{2}$ are considerably more favorable than the estimated distribution of $X_{1}$, the agent still opts for a complete investment in $X_{2}$ despite the fear of misspecification.
Figure 2: Varying $\sigma_{2}$, $m_{1}=0$, $\sigma_{1}=0.1$, $c=0.5$, $\alpha=0.95$, $\beta=0.95$; the red dotted curve connects the optimizers for different $\sigma_{2}$

### v5

@@ 6.2 Two constructions of partially law-invariant coherent risk measures
where $\widehat{\rho}^{\beta}$ is the $\beta$-tail risk measure generated by $\widehat{\rho}$. See Section 13.2 for a proof of this claim.
6.3 Numerical illustrations
In this section, we numerically evaluate the formula from Proposition 6.3 to generate plots and explore how model uncertainty shapes portfolio allocations.
Let $(\Omega,\mathcal{F})=(\mathbb{R}^{2},\mathcal{B}(\mathbb{R}^{2}))$ and $\mathbb{P}$ be multivariate Gaussian with mean $m\in\mathbb{R}^{2}$ and covariance matrix $\Sigma\in\mathbb{R}^{2\times 2}$, where $\Sigma_{ij}=(c+(1-c)\mathds{1}_{\{i=j\}})\sigma_{i}\sigma_{j}$ for $i,j\in\{1,2\}$ (that is, an equicorrelation matrix). Let $X_{1},X_{2}$ be the random variables given by the projections, representing the single-period percentage losses of two financial positions. We interpret $\mathbb{P}$ as the initial estimated bivariate distribution of these losses. Assuming the distribution of $X_{1}$ is accurately specified (see Section 2.2), set $\mathcal{G}=\sigma(X_{1})$. The objective is to evaluate the risk of portfolios of the form $\pi_{1}X_{1}+\pi_{2}X_{2}$ using the risk measure $\rho_{\beta}=\sup_{\mu\in\mathcal{Q}_{\beta}}\mathrm{ES}^{\mu}_{\alpha}$, where $(\pi_{1},\pi_{2})$ lies in the unit simplex of $\mathbb{R}^{2}$. The parameter $\beta\in[0,1)$ represents the decision maker’s confidence in $\mathbb{P}$ for the events outside $\mathcal{G}$. When $\beta=0$, the risk measure $\rho_{0}$ coincides with $\mathrm{ES}_{\alpha}$, reflecting complete confidence in the reference probability model $\mathbb{P}$. As $\beta$ increases, it signals a growing doubt in the model’s accuracy outside of $\mathcal{G}$, leading to more conservative risk evaluations.
In Section 13.3, for $\beta\in[0,1)$, we derive an integral formula for
@@ 6.3 Numerical illustrations
which can be evaluated using numerical integration. By Proposition 6.3,
$\rho_{\beta}\left(\pi_{1}X_{1}+\pi_{2}X_{2}\right)=\min_{x\in\mathbb{R}}\left(x+\frac{1}{1-\alpha}f_{\beta}(\pi_{1},\pi_{2},x)\right),$
which can be computed numerically using a convex program.
Figures 1 and 2 display plots of the computed values of $\rho_{\beta}$ for various portfolio weights, modeling assumptions, and values of $\beta$. The horizontal axes in all panels are given by $\pi_{1}$, the weight allocated to $X_{1}$. Since $\rho_{\beta}(X_{1})=\mathrm{ES}_{\alpha}(X_{1})$ for all $\beta\in[0,1)$, the lines agree when $\pi_{1}=1$.
Figure 1: Varying $\beta$, while fixing $m_{1}=0$, $\sigma_{1}=0.1$, $\sigma_{2}=0.1$, $c=0.5$, $\alpha=0.95$; the red dotted curve connects the optimizers for different $\beta$
Figure 2: Varying $\sigma_{2}$, while fixing $m_{1}=0$, $\sigma_{1}=0.1$, $c=0.5$, $\alpha=0.95$, $\beta=0.95$; the red dotted curve connects the optimizers for different $\sigma_{2}$
Figure 1 illustrates how the risk-minimizing portfolio evolves as $\beta$ increases. When $m_{2}=0$ (left panel), an agent with $\beta=0$ optimally selects an equal allocation between $X_{1}$ and $X_{2}$. This is consistent with the preference for pure diversification as both $X_{1}$ and $X_{2}$ share the same correctly-specified distribution. As $\beta$ increases, the risk-minimizing portfolio shifts monotonically toward a full allocation to $X_{1}$. This shift reflects the agent’s growing concern that the model may be misspecified outside the estimated distribution for $X_{1}$. The center and right panels ($m_{2}=-0.05$ and $-0.1$) represent cases where $X_{2}$ is initially expected to yield a higher return. In these cases, the same shift towards a full investment into $X_{1}$ persists, albeit with reduced intensity as the perceived downside of model uncertainty is partially offset by the favorable mean of $X_{2}$.
Figure 2 considers a fixed $\beta=0.95$ and varying $\sigma_{2}$. As expected, when $\sigma_{2}$ is very small, even the worst-case distributions within the “$\beta$-neighborhood” of the estimated distribution for $X_{2}$ are more favorable than the well-specified distribution of $X_{1}$. Consequently, the agent continues to allocate fully to $X_{2}$, despite the high level of model uncertainty. When $\sigma_{2}$ is large, the risk-minimizing portfolio has an increasing weight allocated to $X_{1}$.